I am a Grandmother trying to help my grandson with his homework. I haven't taken Algebra in over 45 years. Please help me. This is a work problem from a print out so I don't know the text book ISBN.
Tyrone is twice as old as his daughter. In six years Tyrone's age will be three times what his daughter's age was six years ago. How old are they at present?

Answer by feliz1965(151)

Tyrone is twice as old as his daughter. In six years Tyrone's age will be three times what his daughter's age was six years ago. How old are they at present?
Tyrone's age now = 2x
His daughter's age now = x
2x + 6 = 3(x - 6)
2x + 6 = 3x - 18
2x - 3x = -6 - 18
-x = -24
x = -24/-1
x = 24
Tyrone's daughter is 24 years old and he is twice that age or 48 years old.